Proof that System -y 9--x-x has no Periodic trayectory C0ueXrerience a Sacd lera
Answer #1

Given non-linear system is. si =y y = - - (Itatntly Flm, = y; alms)=– n. (1ttently clearez cool is the equillibrium point. Th& .= -3 +in: 22 = -12 - 18 are the complex pair of eigenra with negative real part. & The system is stable and the Critical p